OpenStudy (anonymous):

How much work is required to assemble eight identical point charges, each of magnitude q, at the corners of a cube of side s? Note: Assume a reference level of potential V = 0 at r = . (Use any variable or symbol stated above along with the following as necessary: ke.)

OpenStudy (anonymous):

just sum the voltages and multiply by q: ie each charge must move in against the fields of the charges already in place... so the work for each is q* (V1 +V2 +V3......) the first charge doesn't have to move in against any field: W=0 the second charge has to move in to a distance 's' from the first charge: W= q(kq/s) the third charge has to move in against the fields of the first two charges to a distance of s from one and sqrt(2)*s from the other : W=q( kq/s +kq/(sqrt(2)*s) )  and so on...

OpenStudy (anonymous):

What are the Magnitude and direction of this resulant force?

OpenStudy (anonymous):

The resultant magnitude is the sum of all the individual magnitudes, the direction (if its a perfect cube) is 0, since each particle has an exact opposite to cancel it out. each pair is opposite so cancels, 1+2 cancel, 3+4 cancel etc.